OpenStudy (cutiecomittee123):

The population (in millions) of a certain country can be approximated using the following function; P(x)=40 ( 1.02)^x where x is the number of years after 2000. which of the following calculations will tell you what the population can be expected to reach 200 million?

OpenStudy (zmtgaminghd):

279,378

OpenStudy (cutiecomittee123):

oh well there are options ln(5/1.02) ln(5)/ln(1.02)+2000 ln(5/1.02)+2000 ln(5)/ln(1.02)

OpenStudy (campbell_st):

ok so let P(x) = 200 so \[200 = 40\times (1.02)^x\] divide both sides of the equation by 40 \[5 = (1.02)^x\] take the log of both sides... and apply the log law for powers and you'll get the answer

OpenStudy (campbell_st):

the log law for powers is \[\log(x^a) = a \times \log(x)\]

OpenStudy (cutiecomittee123):

so log(5)=log(1.02)^x ??

OpenStudy (campbell_st):

yes... so now what happens to x, with you use the log law for powers

OpenStudy (cutiecomittee123):

can you show me how that works?

OpenStudy (campbell_st):

ok \[\log(5) = x \times \log(1.02)\] now you can solve for x

OpenStudy (cutiecomittee123):

got it! so just divide log(5) by log(1.02)

OpenStudy (campbell_st):

that's correct

OpenStudy (cutiecomittee123):

so the answer is d, in the options i have above?

OpenStudy (campbell_st):

yes

OpenStudy (cutiecomittee123):

Can you help me with another one?

OpenStudy (cutiecomittee123):

@campbell_st

OpenStudy (campbell_st):

just post a new question, there are lots of people who can help, I'll look at it if I have time

OpenStudy (cutiecomittee123):

Okay